OpenStudy (anonymous):

Using the given zero, find all other zeros of f(x). -2i is a zero of f(x) = x4 - 32x2 - 144

OpenStudy (anonymous):

if \(2i\) is a zero then so is \(-2i\) therefore one factor is \(x+2i\) and another is \(x-2i\) multiply the factors, get \[x^2+4\]

OpenStudy (anonymous):

therefore this sucker factors as \[x^4-32x^2-144=(x^2+4)(\text{something})\]

OpenStudy (anonymous):

2i, 12, -12 2i, 6i, -6i 2i, 6, -6 2i, 12i, -12i

OpenStudy (anonymous):

on the other hand you can solve this directly without knowing any zeros, put \(u=x^2\) and solve \[u^2-32u-144=0\] by factoring

OpenStudy (anonymous):

those are my options

OpenStudy (anonymous):

once you find \(u\) then replace it by \(x^2\) and solve

OpenStudy (anonymous):

you know how to factor this ?

OpenStudy (anonymous):

hint, one of the factors is \(x^2+4\)

OpenStudy (anonymous):

is the answer a?

OpenStudy (anonymous):

probably want me to check?

OpenStudy (anonymous):

yes please :)

OpenStudy (anonymous):

actually i change my mind probably not

OpenStudy (anonymous):

\[x^4-31x-144=0\\ (x^2+4)(x^2+bx+c)=0\]can factor easily how many times does \(4\) go in to \(144\)?

OpenStudy (anonymous):

36

OpenStudy (anonymous):

actually \(-36\)

OpenStudy (anonymous):

so this factors as \[(x^2+4)(x^2-36)=0\]

OpenStudy (anonymous):

final job it to solve \[x^2-36=0\] for \(x\)

OpenStudy (anonymous):

6, -6

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

Oh i see what i did wrong now, thanks!

OpenStudy (anonymous):

yw